# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ 3.2 Formal Modeling and Realization of Training Information
$I_{t}=\langle ot,T_{th},S_{ot},M_{I},t_{1},S_{M_{I}}\rangle$ (3.7)
$S_{ot}(ot,T_{th})=\{\varphi_{t}|\varphi_{t}\text{ is a WFF about }ot\text{ and }T_{th}\}$ (3.8)
$S_{M_{I}}(M_{I},t_{1})=\{\psi_{t}|\psi_{t}\text{ is a WFF about }M_{I}\text{ and }t_{1}\}$ (3.9)
Here, the ontology $ot$ refers to the object containing knowledge or samples; $T_{th}$ is the time at which its state is formed; and $S_{ot}(ot,T_{th})$ is a set of business-oriented semantic assertions or facts about the ontology $M_{I}$ functions similarly to the hardware interface and API software of M, carrying the training information at time $t_{1}$. The carrier state $S_{M_{I}}(M_{I},t_{1})$ is the physical realization of $S_{ot}(ot,T_{th})$ via an enabling mapping, ensuring that the information can be machine-read and entered into M.
@@ 3.3 Formal Modeling and Implementation of Model Learning
Model learning for M involves fusing its original information $I_{o}$ with the training information $I_{t}$, resulting in updated information:
$I_{u}=\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\rangle$ (3.10)
@@
$S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})=\{\varphi_{u_{1}},\varphi_{u_{2}},\varphi_{u_{3}}\}$ (3.11)
$\varphi_{u_{1}}=Eq^{3}(\Phi^{2}(M,t_{2}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.12)
indicates that the input state $S_{M_{I}}(M_{I},t_{1})$, which was learned by M at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of M at $t_{2}$.
$\varphi_{u_{2}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.13)
@@
$\varphi_{u_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Processable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.14)
indicates that when the training information $I_{t}$ is processable by M, it can also produce output information.
@@
$S_{M}(M,t_{2})=\{\psi_{u_{1}},\psi_{u_{2}},\psi_{u_{3}}\}$ (3.15)
3.4 Information Processing and Learning Optimization in Models
@@
$I_{q}=\langle oq,T_{qh},S_{oq},M_{I},t_{2},S_{M_{I}}\rangle$ (3.16)
$S_{oq}(oq,T_{qh})=\{\varphi_{q}|\varphi_{q}\text{ is a WFF about }oq\text{ and }T_{qh}\}$ (3.17)
$S_{M_{I}}(M_{I},t_{2})=\{\psi_{q}|\psi_{q}\text{ is a WFF about }M_{I}\text{ and }t_{2}\}$ (3.18)
Here, the ontology $oq$ represents the user’s intention or device, with $T_{qh}$ denoting the time at which the ontological state is formed. The state $S_{oq}(oq,T_{qh})$ is the formal representation of the user’s needs. The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,T_{qh})$ onto the model’s interface device $M_{I}$.
@@ 3.5 Information Processing, Output, and Online Learning of the Model
$I_{v}=\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\rangle$ (3.19)
@@
$S_{M\cup M_{I}\cup M_{O}}(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3})=\{\varphi_{v_{1}},\varphi_{v_{2}},\varphi_{v_{3}}\}$ (3.20)
$\varphi_{v_{1}}=State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.21)
@@
$\varphi_{v_{2}}=$ $Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land$
$Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.22)
indicates that when the input state $S_{M_{I}}(M_{I},t_{2})$ is received at $t_{2}$, M produces the output state $S_{M_{O}}(M_{O},t_{3})$ at $t_{3}$.
$\varphi_{v_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.23)
@@
The physical realization of this ontological state through an enabling mapping is:
$S_{M}(M,t_{3})=\{\psi_{v_{1}},\psi_{v_{2}},\psi_{v_{3}}\}$ (3.24)
@@
$I_{r}=\langle M\cup M_{I},t_{2},S_{M\cup M_{I}},M_{O},t_{3},S_{M_{O}}\rangle$ (3.25)
@@
$S_{M\cup M_{I}}(M\cup M_{I},t_{2})=\{\varphi_{r_{1}}\}$ (3.26)
@@
$\varphi_{r_{1}}=Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))$ (3.27)
@@ 3.6 Multi-Stage Information Causal Chains
From a systems engineering perspective, the realization of information in machine learning involves more than just five information mappings. In practice, stages such as requirements analysis, design and development, and system integration are also necessary. Specifically, in the four stages of model construction, learning and training, problem input, and answer output, and according to the corollary on information chains in Section 2.5, there is a causal and progressive process that moves from conceptual ideas to mind maps, program code, and finally to physical implementation. In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Figure.1).
Figure 1: Machine Learning Causality Chains with Multiple Information Mappings
As shown in Figure 1, each of these four stages encompasses four layers: conceptual ideas, mind maps, program code, and physical models. Typically, content within the same layer shares the same semantic logic and can thus be uniformly represented by a formal system. Furthermore, due to the transmissive nature of information, the innermost physical model can also be regarded as the result of enabling mappings from outer conceptual ideas or mind maps. This illustrates that the entire system comprises multiple stages, with each stage involving various combinations of information causal chains.
This framework, based on the realization of information and finite automaton states, provides a formalized representation and a comprehensive, causally progressive structure for each phase in the machine learning process. It transcends specific algorithms or models, offering a unified and universal framework, MLT-MF. In the appendix, we also present an example of its application to feedforward neural networks, which validates the correctness and applicability of the framework and supports the verification of the related theorems.
4 Several Important Theorems and Proofs
@@ 4.1 Theorem of Model Interpretability
Interpretability of complex models, such as those in deep learning, has long been one of the key challenges in artificial intelligence. However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability Lipton2018 ; Doshi2017 ; Rubin2019 ; Miller2019 . In fact, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system. Furthermore, it should also be considered in relation to the specific inputs and outputs of the model. Based on MLT-MF, we define the following concepts:
@@
Definition 9 (Model Interpretability) Let the model M, as a finite automaton, operate over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, at which it undergoes assignment, input, output, and state transitions. The state of M at time $t_{i}$ is denoted as $S_{M}(M,t_{i})$, and the state of its input device $M_{I}$ at time $t_{i}$ is $S_{M_{I}}(M_{I},t_{i})$, while the state of the output device $M_{O}$ at time $t_{i+1}$ is $S_{M_{O}}(M_{O},t_{i+1})$. If there exists a formal system $\mathcal{L}$ along with three interpretation relations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$ respectively, such that
$E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$ (4.1)
@@
making $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ the respective enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, then M is said to be interpretable in $\mathcal{L}$ with respect to output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$. If for all $i=1,\cdots,n-1$, the output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in $\mathcal{L}$, then M is called interpretable in $\mathcal{L}$.
@@
Here, $\text{Cn}(E_{M}\cup E_{M_{I}})$ denotes the logical closure of $E_{M}\cup E_{M_{I}}$. Equation 4.1 states that in the formal system $\mathcal{L}$, all formulas in $E_{M_{O}}$ are logical consequences of $E_{M}$ and $E_{M_{I}}$. Moreover, since $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are respectively enabled by $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, the formal system $\mathcal{L}$ and its three interpretations fully capture the logical and causal relationship whereby M produces output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$. This, in essence, offers a concise understanding of model interpretability.
@@
Theorem 4 (Model Interpretability and Information Existence) Let the model M be viewed as a finite automaton operating over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, during which it undergoes assignment, input, output, and state transitions.
@@
(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}(M_{I},t_{i})$ is that there exist $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$ and $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$ representing the information carried by M and its input device $M_{I}$ at time $t_{i}$ respectively, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$. The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$ must hold.
@@
Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$. According to Definition 9, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$, and $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
Let $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$. Then, according to Definition 6,we find there exist information objects $Iu(t_{i})=$ $\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, similarly , then $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),$ $Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Thus, the necessity of the first part of the theorem is established.

### v10

@@ 4.2 Formal Modeling and Realization of Training Information
$It=\left\langle ot,{Tt}_{h},S_{ot},M_{I},t_{1},S_{M_{I}}\right\rangle$ (4.7)
$S_{ot}(ot,{Tt}_{h})=\left\{\varphi t\middle|\varphi t\ is\ a\ WFF\ about\ ot\ and\ {Tt}_{h}\right\}$ (4.8)
$S_{M_{I}}\left(M_{I},t_{1}\right)=\left\{\psi t\middle|\psi t\ is\ a\ WFF\ about\ M_{i}\ and\ t_{1}\right\}$ (4.9)
Here, the ontology $ot$ refers to the object containing knowledge or samples; ${Tt}_{h}$ is the time at which its state is formed; and $S_{ot}(ot,{Tt}_{h})$ is a set of business-oriented semantic assertions or facts about the ontology. $M_{I}$ functions similarly to the hardware interface and API software of $M$, carrying the training information at time $t_{1}$. The carrier state $S_{M_{I}}\left(M_{I},t_{1}\right)$ is the physical realization of $S_{ot}(ot,{Tt}_{h})$ via an enabling mapping, ensuring that the information can be machine-read and entered into $M$.
@@ 4.3 Formal Modeling and Implementation of Model Learning
Model learning for $M$ involves fusing its original information $Io$ with the training information $It$, resulting in updated information:
$Iu=\left\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\right\rangle$ (4.10)
@@
$S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})=\left\{{\varphi u}_{1},{\varphi u}_{2},{\varphi u}_{3}\right\}$ (4.11)
${\varphi u}_{1}={Eq}^{2}(\Phi^{2}\left(M,t_{2}\right),f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right)))\land{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)$ (4.12)
indicates that the input state $S_{M_{I}}\left(M_{I},t_{1}\right)$, which was learned by $M$ at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of $M$ at $t_{2}$.
$\begin{split}{\varphi u}_{2}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\land{Learnable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &({Eq}^{2}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))))\land{State}^{3}\left(M,t_{3},\Phi^{2}(M,t_{3}))\right)\end{split}$ (4.13)
@@
$\begin{split}{\varphi u}_{3}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\land{Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &({Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right)))\land{Output}^{3}(M,t_{3},S_{M_{O}}\left(M_{O},t_{3}\right))))\end{split}$ (4.14)
indicates that when the training information $It$ is processable by $M$, it can also produce output information.
@@
$S_{M}\left(M,t_{2}\right)=\left\{{\psi u}_{1},{\psi u}_{2},{\psi u}_{3}\right\}$ (4.15)
4.4 User Input Information for the Model
@@
$Iq=\left\langle oq,{Tq}_{h},S_{oq},M_{I},t_{2},S_{M_{I}}\right\rangle$ (4.16)
$S_{oq}(oq,{Tq}_{h})=\left\{\varphi q\middle|\varphi q\ is\ a\ WFF\ about\ oq\ and\ {Tq}_{h}\right\}$ (4.17)
$S_{M_{I}}(M_{I},t_{2})=\left\{\psi q\middle|\psi q\ is\ a\ WFF\ about\ M_{I}\ and\ t_{2}\right\}$ (4.18)
Here, the ontology $oq$ represents the user’s intention or device, with ${Tq}_{h}$ denoting the time at which the ontological state is formed. The state $S_{oq}(oq,{Tq}_{h})$ is the formal representation of the user’s needs. The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,{Tq}_{h})$ onto the model’s interface device $M_{I}$.
@@ 4.5 Information Processing, Output, and Online Learning of the Model
$Iv=\left\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\right\rangle$ (4.19)
@@
$S_{M\cup M_{I}\cup M_{O}}\left(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3}\right)=\left\{{\varphi v}_{1},{\varphi v}_{2},{\varphi v}_{3}\right\}$ (4.20)
${\varphi v}_{1}={State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)$ (4.21)
@@
$\begin{split}{\varphi v}_{2}=&{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\land{Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right)))\\ &\land{Output}^{3}(M,t_{3},S_{M_{O}}\left(M_{O},t_{3}\right))\end{split}$ (4.22)
indicates that when the input state $S_{M_{I}}\left(M_{I},t_{2}\right)$ is received at $t_{2}$, $M$ produces the output state $S_{M_{O}}\left(M_{O},t_{3}\right)$ at $t_{3}$.
$\begin{split}{\varphi v}_{3}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land{Learnable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))({Eq}^{2}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))))\\ &\land{State}^{3}\left(M,t_{3},\Phi^{2}(M,t_{3}))\right)\end{split}$ (4.23)
@@
The physical realization of this state through an enabling mapping is:
$S_{M}(M,t_{3})=\left\{{\psi v}_{1},{\psi v}_{2},{\psi v}_{3}\right\}$ (4.24)
@@
$Ir=\left\langle M\cup M_{I},t_{2},S_{M\cup M_{I}},M_{O},t_{3},S_{M_{O}}\right\rangle$ (4.25)
@@
$S_{M\cup M_{I}}\left(M\cup M_{I},t_{2}\right)=\left\{{\varphi r}_{1}\right\}$ (4.26)
@@
${\varphi r}_{1}={Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))$ (4.27)
@@ 4.6 Multi-Stage Information Causal Chains
From systems engineering perspective, the realization of information in machine learning involves more than just five information mappings. In practice, stages such as requirements analysis, design and development, and system integration are also necessary. Specifically, in the four stages of model construction, learning and training, problem input, and answer output, and according to the corollary on information chains in Section 3.3, there is a causal and progressive process that moves from conceptual ideas to mind maps, program code, and finally to physical implementation. In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Fig.2).
Fig.2 presents a comprehensive framework of formal expressions and causal information progression for each stage of the machine learning process, from the perspectives of OIT and finite automaton state implementation. This framework transcends specific algorithms or models, offering a novel, unified, and general framework—MLT-MF—for addressing several key issues in the field of machine learning. In Appendix C, we further provide an example of its application to feedforward neural networks, thereby testing the correctness and applicability of the framework and supporting the verification of related theorems.
Figure 2: Causal Chains of Multiple Information Mappings in Machine Learning.
In machine learning, the four distinct stages—model construction, learning and training, problem input, and answer output—each encompass four levels: conceptual ideas, mind maps, program code, and physical models. Generally, the content within the same level shares a common semantic logic and can thus be represented using a unified formal system. Furthermore, owing to the transmissibility of information, the innermost physical model can be regarded as the result of an enabling mapping from the outer conceptual ideas or mind maps. The output stage can, in turn, be viewed as a reverse causal transmission from the physical model back to conceptual ideas. This demonstrates that the system comprises multiple stages, with each stage containing various combinatorial information causal chains.
5 Several Important Definitions and Theorems
@@ 5.1 Theorem of Model Interpretability
According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model (Glanois et al., 2024), the following concepts are established based on the MLT-MF framework:
@@
Definition 4 (Model Interpretability) Let the model $M$, as a finite automaton, operate over an ordered set of times $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$, at which it undergoes assignment, input, output, and state transitions. The state of $M$ at time $t_{i}$ is denoted as $S_{M}\left(M,t_{i}\right)$, and the state of its input device $M_{I}$ at time $t_{i}$ is $S_{M_{I}}\left(M_{I},t_{i}\right)$, while the state of the output device $M_{O}$ at time $t_{i+1}$ is $S_{M_{O}}\left(M_{O},t_{i+1}\right)$. If there exists a formal system $\mathcal{L}$ along with three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$ respectively, such that
$E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$ (5.1)
@@
making $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ the respective enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, then $M$ is said to be interpretable in $\mathcal{L}$ with respect to output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ under input $S_{M_{I}}\left(M_{I},t_{i}\right)$. If for all $i=1,\cdots,n-1$, the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ under input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in $\mathcal{L}$, then $M$ is called interpretable in $\mathcal{L}$.
@@
Here, $Cn(E_{M}\cup E_{M_{I}})$ denotes the logical closure of $E_{M}\cup E_{M_{I}}$. Equation (5.1) states that in the formal system $\mathcal{L}$, all formulas in $E_{M_{O}}$ are logical consequences of $E_{M}$ and $E_{M_{I}}$.Moreover, since $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are respectively enabled by $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, the formal system $\mathcal{L}$ and its three interpretations fully capture the logical and causal relationship whereby $M$ produces output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ under input $S_{M_{I}}\left(M_{I},t_{i}\right)$. This, in essence, offers a concise understanding of model interpretability.
@@
Theorem 1 (Model Interpretability and Information Existence) Let the model $M$ be viewed as a finite automaton operating over an ordered set of times $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$, during which it undergoes assignment, input, output, and state transitions.
@@
(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is that there exist $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$ and $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$ representing the information carried by $M$ and its input device $M_{I}$ at time $t_{i}$, respectively, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$. The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$ must hold.
@@
Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$. According to Definition 4, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$, and $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
Let
$S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$ $=E_{M},$
$S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$ $=E_{M_{I}},$
$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$ $=E_{M_{O}}$
Then, according to Definition 6, there exist information objects
$Iu(t_{i})$ $=\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle,$
$Iq(t_{i})$ $=\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle,$
$Ir(t_{i+1})$ $=\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$
such that
$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$
$\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$
Thus, the necessity of the first part of the theorem is established.
On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$, $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$, such that in the formal system $\mathcal{L}$,
$S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))$
$\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$